You are testing H0: u=100 against HA: u>100 based on an SRS of 16 observations from a Normal population. The t-statistic is t = 2.13 1.

The degrees for the t statistic are:

A. 15

B. 16.

C. 17

2.

The p-value for the statistic in the previous exercise:

A. falls between 0.05 and 0.10

B. falls between 0.01 and 0.05

C. is less than 0.01
Answer #1

1) degrees of freedom = n -1

d.f = 16 -1

Degrees of freedom = 15

2)

P-value for t = 2.13 , d.f = 15 , right tailed test

p-value = P( t > 2.13) df = 15

p-value = 0.0251

P-value falls between 0.01 and 0.05